# MALADY: Multiclass Active Learning with Auction Dynamics on Graphs

### v2

@@ 2.3 Auction Algorithm with Upper and Lower Bound Volume Constraints
In many scenarios, the exact number of elements in each class is unknown. Thus, in [25], the authors present a upper and lower bound auction algorithm for semi-supervised learning that allows the size of each class to fluctuate between upper and lower bounds. In particular, let $N$ be the number of classes, and suppose that class $j$ must have at least $B_{j}$ members and at most $U_{j}$ members, where $B_{j}\leq U_{j}$, and $\sum_{1}^{N}B_{j}\leq n\leq\sum_{j=1}^{N}U_{j}$, where $n$ is the number of data elements in the data set.
@@
The addition of these bounds in the optimization problem introduces some complexities in the problem. For example, each data element $x$ always aims to obtain the most desirable class $j$, which may result in a deficiency of members in other classes. To solve this problem, the authors of [25] introduce an idea of incentives $\bm{t}$ in the market mechanics. In particular, class $i$ must sell $B_{i}$ memberships and if it is having trouble attracting a certain number of members, it should offer incentives to attract elements to join the class. This results in a competition among the classes, and the classes that are deficient in members are forced to offer competitive incentives to attract the necessary number of data elements. To satisfy the lower bounds, one needs to apply an adaptation of the reverse auction algorithm of Bertsekas and coauthors [43] and have the classes bid on the data elements.
@@
The complementary slackness condition for (LABEL:eq:ulassign) and (14) states that an assignment $\bm{u}$ and dual variables $(\bm{p},\bm{t},\bm{\pi})$ are optimal if and only if
@@
Unfortunately, the last two terms in (LABEL:eq:ulcs) do not have useful $\epsilon$ relaxations. Therefore, the authors of [25] propose a two stage method to solve (LABEL:eq:ulassign). First, one should run the forward auction algorithm, which is Algorithm 4 in the Appendix, to satisfy the upper bound constraints with a complete $\epsilon$-CS matching. The output of this stage is then fed into a lower bound auction, which is Algorithm 5 in the Appendix; the goal of the lower bound auction is to obtain a matching that satisfies the lower bound class size constraints as well. Therefore, at the end, one will obtain a $\epsilon$-CS matching satisfying both the upper and lower bound class constraints. We refer the reader to [25] for a detailed description.
@@ 3.1.1 Notation
We use a graph-based framework to derive our underlying semi-supervised classifier. Let $N$ denote the number of classes, and let the data set $V$ consist of the training elements $X_{l}$ with label information and $X_{u}$, the unlabeled training data elements. We embed our data set into a weighted similarity graph $G(V,W)$ using the similarity functions mentioned in Section 2.1. For class $i$, let $B_{i}$ and $U_{i}$ denote the lower bounds and upper bounds on the class, respectively. If exact class sizes are known, we can set $B_{i}=U_{i}$, and when the class information is not available, we simply set $B_{i}=0$ and $U_{i}=|V|$. Let $l_{x}\in\{1,...,N\}$ denote the label of $x\in X_{l}$, $e_{i}$ denote a vector with $1$ in the $i^{th}$ place and 00 elsewhere.
@@ 3.1.2 Derivation
$E(\bm{\Sigma})=R(\bm{\Sigma})+F(\bm{\Sigma}),$ (17)
where $R(\bm{\Sigma})$ is a regularizing term which ensures the smoothness in the partition $\bm{\Sigma}=(\bm{\Sigma}_{1},...,\bm{\Sigma}_{N})$ into $N$ classes, and $F(\bm{\Sigma})$ is a fidelity term containing information from the labeled training data. We can reformulate (17) by imposing constraints which incorporate the labeled training data and class size information:
$\small B_{i}\leq|\Sigma_{i}|\leq U_{i},\quad X_{i}\subset\Sigma_{i}\quad\text{for all}\quad 1\leq i\leq N,$ (18)
where $U_{i}$ and $B_{i}$ are upper and lower bounds on the class sizes, respectively, and $X_{i}$ is the set of labeled training points with label $i$. Then, to guarantee a notion of smoothness for partitions of the data set $V$, one can use a weighted graph cut as a regularizing term. The weighted graph cut is defined as the following term:
$\text{Cut}(\bm{\Sigma})=\frac{1}{2}\sum_{i=1}^{N}\sum_{x\in\Sigma_{i}}\sum_{y\notin\Sigma_{i}}w(x,y),$ (19)
where the entries $w(x,y)$ describe how strongly the points $x$ and $y$ are connected. Unfortunately, the graph cut problem is NP-hard. Instead, similarly to [25, 44], one can consider the graph heat content (GHC) as a convex relaxations of the graph cut. It is defined as:
$\small\text{GHC}(u,W)=\sum_{i=1}^{N}\sum_{x,y\in V}w(x,y)u_{i}(x)(1-u_{i}(y)),$ (20)
where $u$ is an element of the convex relaxation of the space of $N$-phase partitions of $V$. As long as the matrix $W$, for which $W_{ij}=w(x_{i},x_{j})$, is a positive semi-definite matrix, the above GHC term is a concave term.
@@
$\begin{split}&\min_{u}\;\;\text{GHC}(u,W)+\text{J}_{conc}(u)\\ &\text{subject to}\;\;\sum_{i=1}^{N}u_{i}(x)=1\;\;\;\forall x,\\ &u(x)=e_{l_{x}}\;\;\;\;\forall x\in X_{l},\\ &B_{i}\leq\sum_{x\in X_{u}}u_{i}(x)\leq U_{i}\;\;\;\;\forall i\end{split}$ (21)
over ${u:V\rightarrow\mathbb{R}^{k},u\geq 0}$. In this formulation, $\text{GHC}(u,w)$ is the graph heat content energy in (20), which depends on the graph weights, and $\text{J}_{\text{conc}}(u)$ is any concave function of $u$, including any linear function in $u$.
@@
  • The first term of (LABEL:eq:unsupervised) is a convexified graph cut and its minimizer groups data elements so that those of different classes are as dissimilar as possible.
@@
  • It enables one to incorporate a combination of weighted edge-based, class-based and label-based terms which contain important information about the data set elements. This can improve accuracy.
@@
Overall, each step of the scheme to solve (LABEL:eq:unsupervised) is equivalent to finding the ${(k+1)}^{th}$ iterate $u^{k+1}$ as a solution the following modified assignment problem:
$\small u^{k+1}=$
$\small\operatorname*{arg\,max}_{u:V\rightarrow\mathbb{R}^{k},u\geq 0}\sum_{i=1}^{N}\sum_{x\in X_{u}}(1-{(\nabla\text{J}_{\text{conc}})}_{i}(x)-\sum_{y\notin V^{k}_{i}}w(x,y))u_{i}(x)$ (22)
$\begin{split}&\textrm{s.t.}\small\;\sum_{i=1}^{k}u_{i}(x)=1,u(x)=e_{l_{x}}\;\;\;\forall x\in X_{l},\;\\ &L_{i}\leq\sum_{x\in X_{u}}u_{i}(x)\leq U_{i}\quad\;\forall i,\vspace{-0.125cm}\end{split}$ (23)
At each iteration, a partition of the data can be formed via $V_{i}^{k+1}=\{x\hskip 5.69046pt\textrm{s.t.}\hskip 5.69046pt\operatorname*{arg\,max}_{j}u_{j}^{k+1}(x)=i\}.$ Moreover, the elements $\{a_{i}(x)\}$, where
$a_{i}(x)=(1-{(\nabla\text{J}_{\text{conc}})}_{i}(x)-\sum_{y\notin V^{k}_{i}}w(x,y)),$ (24)
are referred as coefficients of the modified assignment problem. Note that the above scheme is equivalent to the upper and lower bound assignment problem (LABEL:eq:ulassign). Our goal is to solve each step of the scheme in (22)-(LABEL:part2) with an auction-type algorithm which incorporates lower and upper bound class size constraints, and also integrates a novel active learning component which will be described in future sections of the paper.
@@
Another example of $\text{J}_{\text{conc}}$ one can use is the modularity term in [45] : $-\gamma\sum_{i}\sum_{x\in V}k_{x}\lvert u_{i}(x)-\frac{1}{2r}\sum_{y\in V}k_{y}u_{i}(y)\rvert^{2}$, where a total variation optimization problem with this term is shown to be equivalent to modularity optimization, which tries to detect communities in the graph based on their modularity. Here, $k_{x}=\sum_{y}w(x,y)$ is the degree of vertex $x$, $\gamma>0$ is a resolution parameter and $2r=\sum_{x}k_{x}$. Using this term will open new avenues to approach network science and detection of communities.
Lastly, it is clear that any linear term in $u$, such as $\sum_{i=1}^{m}\sum_{x\in V}R_{i}(x)u_{i}(x)$, can be used for $\text{J}_{\text{conc}}$. In particular, one can assign $R_{i}(x)$ to be a class homogeneity term, where $R_{i}(x)$ is the cost of assigning $x$ to class $i$. In the latter case, one should formulate $R_{i}(x)$ so that it is small if $x$ is likely to belong to class $i$, and large otherwise. The terms may be defined using the eigenvectors of the graph Laplacian, or using a fit to an expected value of a variable. Please refer to [46] for a more detailed description of class homogeneity terms.
Input : Vertex set $V$, Weight matrix $W$, initial partition $\bm{V}$, lower bound $\bm{B}$, upper bound $\bm{U}$, time step $\delta t$, number of steps $s$, auction error tolerance $\epsilon_{min}$, epsilon scaling factor $\alpha$, initial epsilon $\epsilon_{0}$.
Result: Final configuration $\bm{V}^{s}$, price vector $\bm{p}$, incentives vector $\bm{t}$
Initialization: Set $\bm{V}^{0}:=\bm{V}$, set $\bar{\epsilon}=\epsilon_{min}/N$;
@@
       Calculate the assignment problem coefficients: $\bm{a}_{i}^{k+1}(x)=(1-{(\nabla\text{J}_{\text{conc}})}_{i}(x)-\sum_{y\notin V^{k}_{i}}w(x,y))\hskip 5.69046pt\forall x,i$;
       Initialize prices $\bm{p}=\bm{0}$, incentives $\bm{t}=\bm{0}$, and $\epsilon=\epsilon_{0}$;
@@
             Run Algorithm 4 (Upper Bound Auction): $(\bm{\Sigma}_{\textrm{out1}},\bm{p}_{\textrm{out1}},\bm{t}_{\textrm{out1}})=\textrm{ UBA }(\epsilon,\bm{B},\bm{U},\bm{a}^{k+1},\bm{p},\bm{t},V)$;
             Run Algorithm 5 (Lower Bound Auction): $(\bm{\Sigma}_{\textrm{out2}},\bm{p}_{\textrm{out2}},\bm{t}_{\textrm{out2}})=\textrm{ LBA }(\epsilon,\bm{B},\bm{U},\bm{a}^{k+1},\bm{p}_{\textrm{out1}},\bm{t}_{\textrm{out1}},\bm{\Sigma}_{\textrm{out1}})$;
@@
return ($\bm{V}^{s},\bm{p},\bm{t}$)
Algorithm 1 Semi-supervised learning (SSL) framework
@@ 3.2 Proposed active learning acquisition function
where $v(x)=\max_{1\leq j\leq m}(a_{j}(x)-p_{j}+t_{j}+\epsilon)$ and $w(x)=\max_{1\leq j\leq m,j\neq j^{*}}(a_{j}(x)-p_{j}+t_{j}+\epsilon).$ Here, $a_{j}(x)$ is the coefficient of the modified assignment problem as defined in (24), $p_{j}$ and $t_{j}$ represent the price and incentives, respectively, for class $i$. As discussed in section 2.3, the term $v(x)$ and $w(x)$ represent the best and second best deal values, respectively, in the market mechanism, offered to $x$ by the end of auction method, and it is important to note that when calculating $\mathcal{M}$, one should use the final price and incentive vector returned from Algorithm 1. Then, the general intuition to motivate our proposed acquisition function is that there is some level of uncertainty as to the optimal class of $x$ when the value of $\mathcal{M}(x)$ is relatively small, and there is little uncertainty as to the optimal class when $\mathcal{M}(x)$ is big.
@@
Figure 2: (a) Ground truth with two initial points, (b-d) Active learning acquisition function value for our proposed acquisition function (26) at different stages of the experiment: (b) Iteration 5, (c) Iteration 50, (d) Iteration 150. Brighter regions of the plot indicate larger acquisition function values and labeled points are marked in red circles, whereas the query point for the current iteration is marked as a red star with a black outline.
@@
Figure 3: The flowchart of our MALADY. Green box: Similarity graph construction using a chosen similarity function (section 2.1); Blue box: Active learning process using proposed acquisition function (section 3.2); Grey box: Auction dynamics algorithm for semi-supervised inference (section 3.1). The result is the partition of the vertex set; Red box: When $|\mathcal{L}|=T$ is met, the accuracy of the proposed method is computed on the unlabeled set.
@@
The acquisition function (26) is a novel uncertainty sampling acquisition function [1] where the optimal price vector $\bm{p}$, optimal incentives vector $\bm{t}$, auction coefficient $\bm{a}$ and epsilon $\epsilon$ are used in each active iteration to determine the query points. Overall, one can think of this acquisition function as prioritizing minimum margin sampling, where the margin is defined as the difference between the best and second best deal values offered to $x$ at the end of the auction algorithm. The strategy is to query a point that has the smallest margin, as a smaller margin indicates greater uncertainty in the decision.
@@
To gain better intuition, we run an experiment on a toy dataset in two dimensions to visualize which points are emphasized by the proposed acquisition function. The dataset consists of six Gaussian clusters where each cluster contains 400 points and has a standard deviation of 0.25, and where the centers lie evenly spaced along the unit circle. To add complexity to the dataset, the classification of the clusters is assigned in an alternating pattern, as illustrated in Figure 2(a). With two initial labeled points, we sequentially query 150 points for labeling via the proposed acquisition function (26). Heatmaps of the acquisition function values for our proposed acquisition function at different stages of the active learning process are shown in Figure 2(b-d). Note that brighter regions in these plots represent larger acquisition function values, helping us to visualize where the largest acquisition function values are concentrated. In our proposed active learning framework, the query point that is selected at each corresponding iteration would be the maximizer of the proposed acquisition function, representing the “brightest” point in the corresponding heatmap.
@@
The main observation from this toy experiment is that the proposed acquisition function prioritizes data points along the boundaries between the clusters. Specifically, the points at the boundaries with more overlap between the clusters have the highest acquisition function values. In particular, we notice that a region that bisects the circle is emphasized at each iteration, where the orientation of that bisecting line depends on the relative number of labeled points in each cluster. A key observation, however, is that these bisecting regions cut through the low-density regions between clusters. Thus, as the active learning process proceeds, the acquisition function focuses on each of the regions between the different clusters. These regions correspond to decision boundaries between the different classes. Overall, successful active learning acquisition functions need to efficiently prioritize these decision boundary regions (often referred to as “exploitation” of decision). The results of this toy experiment suggest that the proposed acquisition function can efficiently prioritize decision boundaries modeled by regions between clusters in the corresponding similarity graph structure.
@@
Input : Vertex set $V$, weight matrix $W$, initial partition $\bm{V}$, lower bound $\bm{B}$, upper bound $\bm{U}$, time step $\delta t$, number of steps $s$, auction error tolerance $\epsilon_{min}$, epsilon scaling factor $\alpha$, initial epsilon $\epsilon_{0}$, initial labeled set $\mathcal{L}_{0}$ , acquisition function $\mathcal{A}(x)$, and budget $T$.
@@
      Run Algorithm 1 (Semi-supervised classification method):  $(\bm{V}^{s},\bm{p},\bm{t})=\text{ SSL }(W,\bm{V},B,U,\delta t,s,\epsilon_{min})$;
      Pick $x^{\ast}\in\mathcal{U}=V\setminus\mathcal{L}$ which has highest acquisition value $\mathcal{A}(x)$;
@@
Run Algorithm 1 with labeled set $\mathcal{L}$ (Semi-supervised classification method):  $(\bm{V}^{s},\bm{p},\bm{t})=\text{ SSL }(W,\bm{V},B,U,\delta t,s,\epsilon_{min})$;
@@
return Evaluation of algorithm using chosen evaluation metric

### v3

@@ 2.3 Auction Algorithm with Upper and Lower Bound Volume Constraints
In many scenarios, the exact number of elements in each class is unknown. Thus, in [26], the authors present an upper and lower bound auction algorithm for semi-supervised learning that allows the size of each class to fluctuate between upper and lower bounds. In particular, let $K$ be the number of classes, and suppose that class $j$ must have at least $B_{j}$ members and at most $U_{j}$ members, where $B_{j}\leq U_{j}$, and $\sum_{1}^{K}B_{j}\leq N\leq\sum_{j=1}^{K}U_{j}$, where $N$ is the number of data elements in the data set.
@@
The addition of these bounds in the optimization problem introduces some complexities in the problem. For example, each data element $x$ always aims to obtain the most desirable class $j$, which may result in a deficiency of members in other classes. To solve this problem, the authors of [26] introduce an idea of incentives $\bm{t}$ in the market mechanics. In particular, class $i$ must sell $B_{i}$ memberships, and if it is having trouble attracting a certain number of members, it should offer incentives to attract elements to join the class. This results in competition among the classes, and the classes that are deficient in members are forced to offer competitive incentives to attract the necessary number of data elements. To satisfy the lower bounds, one needs to apply an adaptation of the reverse auction algorithm of Bertsekas and coauthors [51] and have the classes bid on the data elements.
@@
The complementary slackness condition for (LABEL:eq:ulassign) and (14) states that $\bm{u}$ and $(\bm{p},\bm{t},\bm{\pi})$ are optimal if and only if
@@
Unfortunately, the last two terms in (LABEL:eq:ulcs) do not have useful $\epsilon$ relaxations. Thus, [26] proposes a two-stage method to solve (LABEL:eq:ulassign). First, one should run the forward auction algorithm, Algorithm 4 in the Appendix, to satisfy the upper bound constraints with a complete $\epsilon$-CS matching. The output of this stage is then fed into a lower bound auction, Algorithm 3 in the Appendix; the goal of the lower bound auction is to obtain a matching that satisfies the lower bound class size constraints as well. Thus, at the end, one will obtain a $\epsilon$-CS matching satisfying both the upper and lower bound class constraints. We refer the reader to [26] for a detailed description.
@@ 3.1.1 Notation
We use a graph-based framework to derive our underlying semi-supervised classifier. Let $K$ denote the number of classes, and let the data set $\mathcal{X}$ consist of the training elements $\mathcal{L}$ with label information and $\mathcal{U}$, the unlabeled training data elements. We embed our data set into a weighted similarity graph $G(\mathcal{X},W)$ using the similarity functions mentioned in Section 2.1. For class $i$, let $B_{i}$ and $U_{i}$ denote the lower bounds and upper bounds of the class, respectively. If exact class sizes are known, we can set $B_{i}=U_{i}$, and when the class information is not available, we simply set $B_{i}=0$ and $U_{i}=|\mathcal{X}|$. Let $l_{x}\in\{1,...,K\}$ denote the label of $x\in\mathcal{L}$, $e_{i}$ denote a vector with $1$ in the $i^{th}$ place and 00 elsewhere.
@@ 3.1.2 Derivation
$E(\bm{\mathcal{X}})=R(\bm{\mathcal{X}})+F(\bm{\mathcal{X}}),$ (17)
where $R(\bm{\mathcal{X}})$ is a regularizing term which ensures the smoothness in the partition $\bm{\mathcal{X}}=({\mathcal{X}}_{1},...,{\mathcal{X}}_{K})$ into $K$ classes, and $F(\bm{\mathcal{X}})$ is a fidelity term containing information from the labeled training data. We can reformulate (17) by imposing constraints that incorporate the labeled training data and class size information:
$\small B_{i}\leq|\mathcal{X}_{i}|\leq U_{i},\quad\mathcal{L}_{i}\subset\mathcal{X}_{i}\quad\text{for all}\quad 1\leq i\leq K,$ (18)
where $U_{i}$ and $B_{i}$ are upper and lower bounds on the class sizes, respectively, and $\mathcal{L}_{i}$ is the set of labeled training points with label $i$. Then, to guarantee a notion of smoothness for partitions of the data set $\mathcal{X}$, one can use a weighted graph cut as a regularizing term. The weighted graph cut is defined as the following term:
$\text{Cut}(\bm{\mathcal{X}})=\frac{1}{2}\sum_{i=1}^{K}\sum_{x\in\mathcal{X}_{i}}\sum_{y\notin\mathcal{X}_{i}}w(x,y),$ (19)
where the entries $w(x,y)$ describe how strongly the points $x$ and $y$ are connected. Unfortunately, the graph cut problem is NP-hard. Instead, similarly to [26, 52], one can consider the graph heat content (GHC) as a convex relaxation of the graph cut. It is defined as:
$\small\text{GHC}(\bm{u},W)=\sum_{i=1}^{K}\sum_{x,y\in\mathcal{X}}w(x,y)u_{i}(x)(1-u_{i}(y)),$ (20)
where $\bm{u}:\mathcal{X}\to\mathbb{R}^{K}$ is an element of the convex relaxation of the space of $K$-phase partitions of $\mathcal{X}$. For each $x\in\mathcal{X},u(x)=(u_{1}(x),\cdots,u_{K}(x))\in\mathbb{R}^{K}$, where $u_{i}(x)$ is the $i^{th}$ entry of vector $u(x)$. As long as the matrix $W$, for which $W_{ij}=w(x_{i},x_{j})$, is a positive semi-definite matrix, the above GHC term is a concave term.
@@
$\begin{split}&\min_{\bm{u}}\;\;\text{GHC}(\bm{u},W)+\text{J}_{conc}(\bm{u})\\ &\text{subject to}\;\;\sum_{i=1}^{K}u_{i}(x)=1\;\;\;\forall x,\\ &u(x)=e_{l_{x}}\;\;\;\;\forall x\in\mathcal{L},\\ &B_{i}\leq\sum_{x\in\mathcal{U}}u_{i}(x)\leq U_{i}\;\;\;\;\forall i\end{split}$ (21)
over ${\bm{u}:\mathcal{X}\rightarrow\mathbb{R}^{K},\bm{u}\geq 0}$. In this formulation, $\text{GHC}(\bm{u},W)$ is the graph heat content energy in (20), which depends on the graph weights, and $\text{J}_{\text{conc}}(\bm{u})$ is any concave function of $u$, including any linear function in $\bm{u}$.
The $\text{J}_{\text{conc}}(\bm{u})$ term is a term that allows one to consider a wide range of optimization problems within our framework. Without this term, the problem would only involve minimizing the graph heat content subject to certain constraints. The $\text{J}_{\text{conc}}(\bm{u})$ term and some examples of it one can use are described at the end of Section IIIA.
@@
  • The first term of (LABEL:eq:unsupervised) is a convexified graph cut, and its minimizer groups data elements so that those of different classes are as dissimilar as possible.
@@
  • It enables one to incorporate a combination of weighted edge-based, class-based, and label-based terms that contain important information about the data set elements. This can improve accuracy.
@@
Overall, each step of the scheme to solve (LABEL:eq:unsupervised) is equivalent to finding the ${(k+1)}^{th}$ iterate $\bm{u}^{k+1}$ as a solution of the following modified assignment problem:
$\small\operatorname*{arg\,max}_{\bm{u}:\mathcal{X}\rightarrow\mathbb{R}^{K},\bm{u}\geq 0}\sum_{i=1}^{K}\sum_{x\in\mathcal{U}}(1-{(\nabla\text{J}_{\text{conc}})}_{i}(x)-\sum_{y\notin\mathcal{X}_{i}^{k}}w(x,y))u_{i}(x)$ (22)
$\begin{split}&\textrm{s.t.}\small\;\sum_{i=1}^{K}u_{i}(x)=1,u(x)=e_{l_{x}}\;\;\;\forall x\in\mathcal{L},\;\\ &L_{i}\leq\sum_{x\in\mathcal{U}}u_{i}(x)\leq U_{i}\quad\;\forall i,\vspace{-0.125cm}\end{split}$ (23)
At each iteration, a partition of the data can be formed via $\mathcal{X}_{i}^{k+1}=\{x\hskip 5.69046pt\textrm{s.t.}\hskip 5.69046pt\operatorname*{arg\,max}_{j}u_{j}^{k+1}(x)=i\}.$ Moreover, the elements $\{a_{i}(x)\}$, where
$a_{i}(x)=(1-{(\nabla\text{J}_{\text{conc}})}_{i}(x)-\sum_{y\notin\mathcal{X}_{i}^{k}}w(x,y)),$ (24)
are referred to as coefficients of the modified assignment problem. Note that the above scheme is equivalent to the upper and lower bound assignment problem (LABEL:eq:ulassign). Our goal is to solve each step of the scheme in (22)-(LABEL:part2) with an auction-type algorithm that incorporates lower and upper bound class size constraints and also integrates a novel active learning component which will be described in future sections of the paper.
@@
Another example of $\text{J}_{\text{conc}}$ one can use is the modularity term in [53] : $-\gamma\sum_{i}\sum_{x\in\mathcal{X}}k_{x}\lvert u_{i}(x)-\frac{1}{2r}\sum_{y\in V}k_{y}u_{i}(y)\rvert^{2}$, where a total variation optimization problem with this term is shown to be equivalent to modularity optimization, which tries to detect communities in the graph based on their modularity. Here, $k_{x}=\sum_{y}w(x,y)$ is the degree of vertex $x$, $\gamma>0$ is a resolution parameter, and $2r=\sum_{x}k_{x}$. Using this term will open new avenues to approach network science and detection of communities.
Lastly, it is clear that any linear term in $u$, such as $\sum_{i=1}^{m}\sum_{x\in\mathcal{X}}R_{i}(x)u_{i}(x)$, can be used for $\text{J}_{\text{conc}}$. In particular, one can assign $R_{i}(x)$ to be a class homogeneity term, where $R_{i}(x)$ is the cost of assigning $x$ to class $i$. In the latter case, one should formulate $R_{i}(x)$ so that it is small if $x$ is likely to belong to class $i$ and large otherwise. The terms may be defined using the eigenvectors of the graph Laplacian or using a fit to an expected value of a variable. Please refer to [54] for a more detailed description of class homogeneity terms.
Input : Vertex set $\mathcal{X}$, Weight matrix $W$, initial partition $\bm{\mathcal{X}}$, lower bound $\bm{B}$, upper bound $\bm{U}$, time step $\delta t$, number of steps $s$, auction error tolerance $\epsilon_{min}$, epsilon scaling factor $\alpha$, initial epsilon $\epsilon_{0}$.
Result: Final configuration $\bm{\mathcal{X}}^{s}$, price vector $\bm{p}$, incentives vector $\bm{t}$
Initialization: Set $\bm{\mathcal{X}}^{0}:=\bm{\mathcal{X}}$, set $\bar{\epsilon}=\epsilon_{min}/N$;
@@
       Calculate assignment problem coefficients: $\bm{a}_{i}^{k+1}(x)=(1-{(\nabla\text{J}_{\text{conc}})}_{i}(x)-\sum_{y\notin V^{k}_{i}}w(x,y))$;
       Initialize prices $\bm{p}=\bm{0}$, incentives $\bm{t}=\bm{0}$, $\epsilon=\epsilon_{0}$;
@@
             Run Algorithm 4 (Upper Bound Auction): $(\bm{\mathcal{X}}_{\textrm{out1}},\bm{p}_{\textrm{out1}},\bm{t}_{\textrm{out1}})=\textrm{ UBA }(\epsilon,\bm{B},\bm{U},\bm{a}^{k+1},\bm{p},\bm{t},\mathcal{X})$;
             Run Algorithm 3 (Lower Bound Auction): $(\bm{\mathcal{X}}_{\textrm{out2}},\bm{p}_{\textrm{out2}},\bm{t}_{\textrm{out2}})=\textrm{ LBA }(\epsilon,\bm{B},\bm{U},\bm{a}^{k+1},\bm{p}_{\textrm{out1}},\bm{t}_{\textrm{out1}},\bm{\mathcal{X}}_{\textrm{out1}})$;
@@
return ($\bm{\mathcal{X}}^{s},\bm{p},\bm{t}$)
Algorithm 1 Auction Dynamics Semi-supervised learning (SSL) framework
@@ 3.2 Proposed active learning acquisition function
where $v(x)=\max_{1\leq i\leq K}(a_{i}(x)-p_{i}+t_{i}+\epsilon)$ and $w(x)=\max_{1\leq i\leq K,i\neq i^{*}}(a_{i}(x)-p_{i}+t_{i}+\epsilon).$ Here, $a_{i}(x)$ is the coefficient of the modified assignment problem as defined in (24), $p_{i}$ and $t_{i}$ represent the price and incentives, respectively, for class $i$, and $i^{*}=\arg\max_{1\leq i\leq K}(a_{i}(x)-p_{i}+t_{i}+\epsilon)$. As discussed in section 2.3, the term $v(x)$ and $w(x)$ represent the best and second best deal values, respectively, in the market mechanism, offered to $x$ by the end of auction method, and it is important to note that when calculating $\mathcal{M}$, one should use the final price and incentive vector returned from Algorithm 1. Then, the general intuition to motivate our proposed acquisition function is that there is some level of uncertainty as to the optimal class of $x$ when the value of $\mathcal{M}(x)$ is relatively small, and there is little uncertainty as to the optimal class when $\mathcal{M}(x)$ is big.
@@
Figure 2: (a) Ground truth with six initial points, (b) Acquisition function value for our proposed acquisition function (26) at Iteration 100, (c) Acquisition function value of uncertainty sampling [1] with Laplace learning [55] at Iteration 100. Brighter regions of the plot indicate larger acquisition function values, and labeled points are marked in red circles; the query point for the current iteration is marked as a red star with a black outline. Notice that our acquisition function (b) focuses on all decision boundaries between oppositely labeled clusters, whereas standard uncertainty sampling (c) only focuses on a subset of these boundaries.
@@
Figure 3: The flowchart of our MALADY. Green box: Similarity graph construction using a chosen similarity function (section 2.1); Blue box: Active learning process using proposed acquisition function (section 3.2); Grey box: Auction dynamics method for semi-supervised inference (section 3.1). The result is the partition of the vertex set; Red box: When $|\mathcal{L}|=T$, the accuracy of the proposed method is computed.
@@
The acquisition function (26) is a novel uncertainty sampling acquisition function [1] where the optimal price vector $\bm{p}$, optimal incentives vector $\bm{t}$, auction coefficient $\bm{a}$ and epsilon $\epsilon$ are used in each active iteration to determine the query points. Overall, one can think of this acquisition function as prioritizing minimum margin sampling, where the margin is defined as the difference between the best and second-best deal values offered to $x$ at the end of the auction algorithm. The strategy is to query a point that has the smallest margin, as a smaller margin indicates greater uncertainty in the decision.
@@
To gain better intuition, we run an experiment on a toy dataset in two dimensions to visualize which points are emphasized by the proposed acquisition function. The dataset consists of six Gaussian clusters, where each cluster contains 400 points and has a standard deviation of 0.25, and where the centers lie evenly spaced along the unit circle. The classification of the clusters is then assigned in an alternating pattern, as illustrated in Figure 2(a). With six initially labeled points, we sequentially query 100 points for labeling via the proposed acquisition function (26). Heatmaps of the values for our proposed acquisition function 26 and uncertainty sampling [1] with a Laplace learning [55] classifier are shown in panels (b) and (c), respectively. Note that brighter regions in these plots represent larger acquisition function values, helping us to visualize where the largest acquisition function values are concentrated. In our proposed active learning framework, the query point that is selected at each corresponding iteration would be the maximizer of the proposed acquisition function, representing the “brightest” point in the corresponding heatmap.
@@
The main observation from this experiment is that the proposed acquisition function accurately prioritizes data points along all boundaries between oppositely labeled clusters. Specifically, points at boundaries with more overlap between clusters have the highest acquisition function values. This contrasts with a simple uncertainty sampling [1] acquisition function with the Laplace learning classifier [56], which in Figure 2 (c) demonstrates that only a subset of the important boundaries between oppositely labeled clusters are emphasized in sampling. Overall, successful active learning acquisition functions need to efficiently prioritize these decision boundary regions (referred to as “exploitation” in active learning [4]). The results of this experiment suggest that the proposed acquisition function can efficiently prioritize decision boundaries modeled by regions between clusters in the corresponding similarity graph structure.
@@
Input : Vertex set ${\mathcal{X}}$, weight matrix $W$, initial partition $\bm{\mathcal{X}}$, lower bound $\bm{B}$, upper bound $\bm{U}$, time step $\delta t$, number of steps $s$, auction error tolerance $\epsilon_{min}$, epsilon scaling factor $\alpha$, initial epsilon $\epsilon_{0}$, initial labeled set $\mathcal{L}_{0}$ , acquisition function $\mathcal{A}(x)$, and budget $T$.
@@
      Run Algorithm 1 (Semi-supervised method):  $(\bm{\mathcal{X}}^{s},\bm{p},\bm{t})=\text{ SSL }(W,\bm{\mathcal{X}},\mathcal{\mathcal{X}},B,U,\delta t,s,\epsilon_{min})$;
      Pick $x^{\ast}\in\mathcal{U}=\mathcal{X}\setminus\mathcal{L}$ which has highest acquisition value $\mathcal{A}(x)$;
@@
Run Algorithm 1 with labeled set $\mathcal{L}$ (Semi-supervised classification method):  $(\bm{\mathcal{X}}^{s},\bm{p},\bm{t})=\text{ SSL }(W,\bm{\mathcal{X}},\mathcal{X},B,U,\delta t,s,\epsilon_{min})$;
@@
return Evaluation of algorithm